# TimeBridge: Non-Stationarity Matters for Long-term Time Series Forecasting

### v3

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Multivariate time series forecasting aims to predict future changes based on historical observations of time series data, which holds significant applications in fields such as financial investment (Sezer et al., 2020), weather forecasting (Karevan & Suykens, 2020), and traffic flow prediction (Shu et al., 2021). However, the inherent non-stationarity of time series (Kim et al., 2022), characterized by short-term fluctuations and long-term trends, introduces challenges such as spurious regressions, making time series forecasting a particularly complex task.
@@ 2 Related Works
Dependency Modeling focuses on designing methods to capture the relationships within multivariate time series, which can be classified into Channel Independent (CI) and Channel Dependent (CD) methods. CI methods (Zeng et al., 2023; Das et al., 2023; Nie et al., 2023; Dai et al., 2024; Lin et al., 2024) rely exclusively on the historical values of each individual channel for prediction, deliberately avoiding cross-channel interactions. This strategy not only stabilizes the training process but also excels at capturing rapid temporal dynamics unique to each variate. In contrast, CD methods (Wu et al., 2021; Zhou et al., 2022; Wu et al., 2023; Zhang & Yan, 2023; Liu et al., 2024a) leverage the interrelationships between variates for modeling. While these methods utilize more information, they struggle with spurious regressions when modeling short-term dependencies, failing to capture rapid changes effectively.
@@ 4 Experiment
Baselines. For long-term forecasting, we select a diverse set of state-of-the-art baselines representative of recent advancements in time series forecasting, including the Transformer-based DeformableTST (Luo & Wang, 2024), the CNN-based ModernTCN (Donghao & Xue, 2024), the MLP-based TimeMixer (Wang et al., 2024b), as well as other competitive methods such as iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), FEDformer (Zhou et al., 2022), MICN (Wang et al., 2022), TimesNet (Wu et al., 2023), and DLinear (Zeng et al., 2023). For short-term forecasting, we add a well-performing baseline SCINet (Liu et al., 2022a). For financial forecasting, we also incorporate the momentum strategy CSM (Jegadeesh & Titman, 1993) and the reversal strategy BLSW (Poterba & Summers, 1988), along with two classic deep learning models, LSTM (Hochreiter & Schmidhuber, 1997) and Transformer (Vaswani et al., 2017a), to provide a comprehensive evaluation.
Implementation Details. All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014) with a learning rate selected from $\{1e\text{-}3,1e\text{-}4,5e\text{-}4\}$. The number of patches $N$ is set to 30. We adopt a hybrid MAE loss that operates in both the time and frequency domains for stable training (Wang et al., 2024a). For additional details on hyperparameter settings and loss function, please refer to the Appendix D.
@@ 4.1 Long-term Forecasting
Setups. We conduct long-term forecasting experiments on several widely-used real-world datasets, including the Electricity Transformer Temperature (ETT) dataset with its four subsets (ETTh1, ETTh2, ETTm1, ETTm2) (Wu et al., 2021), as well as Weather, Electricity, Traffic, Solar (Liu et al., 2024a) and Climate (Godahewa et al., 2021). These datasets exhibit strong non-stationary characteristics, detailed in Appendix C. Following previous works (Zhou et al., 2021; Wu et al., 2021), we use Mean Square Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. We set the input length $I$ to 720 for our method. For other baselines, we adopt the setting that searches for the optimal input length $I$ and other hyperparameters. Details of the metric and the searching process can be found in Sec. B.1 and Sec. E.1.
Results. As shown in Tab. 1, TimeBridge consistently achieves the best overall performance. Notably, the large-scale Climate dataset, with its 1763 channels, presents substantial challenges due to its high dimensionality and intricate temporal dependencies. TimeBridge adeptly balances non-stationarity and dependency modeling, achieving consistently strong predictive performance. Quantitatively, compared to state-of-the-art methods—Transformer-based DeformableTST (Luo & Wang, 2024), CNN-based ModernTCN (Donghao & Xue, 2024), and MLP-based TimeMixer (Wang et al., 2024b)—TimeBridge reduces MSE and MAE by $1.85\%/2.49\%$, $5.56\%/4.12\%$, and $13.66\%/7.58\%$, respectively.
@@
TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear
@@
ETTm1 0.344 0.379 0.362 0.391 0.348 0.383 0.355 0.380 0.353 0.382 0.420 0.435 0.382 0.422 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 0.246 0.310 0.269 0.329 0.257 0.319 0.257 0.318 0.256 0.317 0.518 0.501 0.292 0.343 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 0.397 0.424 0.439 0.448 0.404 0.423 0.427 0.441 0.413 0.434 0.440 0.463 0.428 0.454 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 0.341 0.382 0.374 0.406 0.328 0.377 0.349 0.397 0.324 0.381 0.809 0.658 0.388 0.434 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 0.219 0.249 0.233 0.271 0.222 0.262 0.226 0.264 0.226 0.264 0.228 0.287 0.305 0.287 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 0.149 0.245 0.164 0.261 0.161 0.261 0.185 0.284 0.159 0.253 0.181 0.279 0.205 0.315 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 0.360 0.255 0.397 0.282 0.391 0.278 0.409 0.279 0.391 0.264 0.523 0.284 0.573 0.347 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 0.181 0.239 0.200 0.260 0.185 0.254 0.193 0.252 0.194 0.245 0.191 0.242 0.243 0.350 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
Climate 1.057 0.494 1.060 0.496 1.060 0.496 1.414 0.606 1.387 0.562 1.170 0.550 1.378 0.662 1.155 0.542 1.166 0.570 1.295 0.569 1.134 0.541
@@ 4.2 Short-term Forecasting
Setups. For short-term forecasting, we conduct experiments on the PeMS datasets (Wang et al., 2024b), which capture complex spatiotemporal correlations among multiple variates across city-wide traffic networks. We use mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE) as evaluation metrics. The input length $I$ is set to 96 and the output length $O$ to 12 for all baselines. Details of datasets and metrics are in Appendix C and Sec. B.2.
@@ 4.3 Financial Forecasting
Setups. We conduct experiments on both the U.S. and Chinese stock markets, including the S&P 500 and CSI 500 indices. Stock price movements are influenced by various factors such as economic indicators, market sentiment, geopolitical events, and company-specific news, leading to high non-stationarity. We predict next-day returns using historical data and generate investment portfolios with a buy-hold-sell strategy (Sanderson & Lumpkin-Sowers, 2018). At day $t+1$ open, traders sell day $t$ stocks and buy top-ranked ones based on predicted returns. Following previous work (Lin et al., 2021), we evaluate performance using Annual Return Ratio (ARR), Annual Volatility (AVol), Maximum Drawdown (MDD), Annual Sharpe Ratio (ASR), Calmar Ratio (CR), and Information Ratio (IR). Details of datasets and metrics are in Appendix C and Sec. B.3.
@@ 5 Ablation Studies
Ablation on Integrated and Cointegrated Attention impact and order. We conduct the following experiments: ① Integrated Attention only, ② Cointegrated Attention only, ③ Integrated Attention followed by Cointegrated Attention, and ④ Cointegrated Attention followed by Integrated Attention, with patch downsampling replaced by upsampling in this case. The results in Tab. 5 show that both ① and ② underperform compared to ③, indicating that both components are beneficial. Additionally, ④ shows the weakest performance, possibly because modeling long-term cointegrated relationships first leads to a loss of important short-term temporal features.
@@ A.2 Cointegration and EG Test
Appendix B Metrics
B.1 Long-term Forecasting
@@
B.2 Short-term Forecasting
@@ Appendix C Datasets
  6. (6)
    Climate dataset (Godahewa et al., 2021) records regional precipitation data across 1,763 stations in Australia.
@@
Appendix D Implementation Details
All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014). The batch size is set to 16 for the Electricity and Traffic datasets, and 32 for all other datasets. All models are trained for 10 epochs. Tab. 8 provides detailed hyperparameter settings for each dataset. For the four ETT datasets, the relatively small number of channels results in less pronounced long-term cointegration relationships, as evidenced by the low EG test results in Tab. 7. Therefore, we focus on modeling short-term intra-variate variations only. To improve training stability, we adopt a hybrid MAE loss that operates in both the time and frequency domains. Compared to MSE, which amplifies differences, MAE provides a more stable optimization process. Additionally, the frequency-domain loss mitigates label autocorrelation, making training more effective. The loss function is defined as:
@@ Appendix D Implementation Details
where FFT denotes the Fast Fourier Transform. $\alpha$ is the hyperparameter.
@@
ETTh1 $2$ 00 $30$ $16$ $1e$-$4$ $128$ $128$ 0.35
ETTh2 $2$ 00 $12$ $12$ $1e$-$4$ $128$ $128$ 0.8
ETTm1 $2$ 00 $15$ $6$ $2e$-$4$ $64$ $128$ 0.8
ETTm2 $2$ 00 $30$ $12$ $2e$-$4$ $32$ $64$ 0.35
@@
Appendix E Full Results
E.1 Main Experiments

### v4

@@ Abstract
Time Series Forecasting, Non-stationary, ICML
@@ 1 Introduction
Multivariate time series forecasting aims to predict future changes based on historical observations of time series data, which holds significant applications in fields such as financial investment (Sezer et al., 2020), weather forecasting (Karevan & Suykens, 2020), and traffic flow prediction (Shu et al., 2021; Miao et al., 2024b). However, the inherent non-stationarity of time series (Kim et al., 2022), characterized by short-term fluctuations and long-term trends, introduces challenges such as spurious regressions, making time series forecasting a particularly complex task.
@@ 2 Related Works
Dependency Modeling focuses on designing methods to capture the relationships within multivariate time series, which can be classified into Channel Independent (CI) and Channel Dependent (CD) methods. CI methods (Zeng et al., 2023; Das et al., 2023; Nie et al., 2023; Dai et al., 2024; Lin et al., 2024; Miao et al., 2025) rely exclusively on the historical values of each individual channel for prediction, deliberately avoiding cross-channel interactions. This strategy not only stabilizes the training process but also excels at capturing rapid temporal dynamics unique to each variate. In contrast, CD methods (Wu et al., 2021; Zhou et al., 2022; Wu et al., 2023; Zhang & Yan, 2023; Liu et al., 2024a; Hu et al., 2025b, a) leverage the interrelationships between variates for modeling. While these methods utilize more information, they struggle with spurious regressions when modeling short-term dependencies, failing to capture rapid changes effectively.
@@ 4 Experiment
Baselines. For long-term forecasting, we select a diverse set of state-of-the-art baselines representative of recent advancements in time series forecasting, including the Transformer-based DeformableTST (Luo & Wang, 2024), the CNN-based ModernTCN (Donghao & Xue, 2024), the MLP-based TimeMixer (Wang et al., 2024b), as well as other competitive methods such as iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), Leddam (Yu et al., 2024), MICN (Wang et al., 2022), TimesNet (Wu et al., 2023), and DLinear (Zeng et al., 2023). For short-term forecasting, we add two well-performing baselines SCINet (Liu et al., 2022a) and DUET (Qiu et al., 2025). For financial forecasting, we also incorporate the momentum strategy CSM (Jegadeesh & Titman, 1993) and the reversal strategy BLSW (Poterba & Summers, 1988), along with two classic deep learning models, LSTM (Hochreiter & Schmidhuber, 1997) and Transformer (Vaswani et al., 2017a), to provide a comprehensive evaluation.
Implementation Details. All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014) with a learning rate selected from $\{1e\text{-}3,1e\text{-}4,5e\text{-}4\}$. The number of patches $N$ is set accordingly to different datasets. We adopt a hybrid MAE loss that operates in both the time and frequency domains for stable training (Wang et al., 2024a). For additional details on hyperparameter settings and loss function, please refer to the Appendix E.
@@ 4.1 Long-term Forecasting
Setups. We conduct long-term forecasting experiments on several widely-used real-world datasets, including the Electricity Transformer Temperature (ETT) dataset with its four subsets (ETTh1, ETTh2, ETTm1, ETTm2) (Wu et al., 2021; Miao et al., 2024a), as well as Weather, Electricity, Traffic, and Solar (Liu et al., 2025a, b). These datasets exhibit strong non-stationary characteristics, detailed in Appendix D. Following previous works (Zhou et al., 2021; Wu et al., 2021), we use Mean Square Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. We set the input length $I$ to 720 for our method. For other baselines, we adopt the setting that searches for the optimal input length $I$ and other hyperparameters. Details of the metric and the searching process can be found in Sec. C.1 and Sec. F.1.
Results. As shown in Tab. 1, TimeBridge consistently achieves the best overall performance. Notably, the large-scale Traffic dataset, with its 862 channels, presents substantial challenges due to its high dimensionality and intricate temporal dependencies. TimeBridge adeptly balances non-stationarity and dependency modeling, achieving consistently strong predictive performance. Quantitatively, compared to state-of-the-art methods—Transformer-based DeformableTST (Luo & Wang, 2024), CNN-based ModernTCN (Donghao & Xue, 2024), and MLP-based TimeMixer (Wang et al., 2024b)—TimeBridge reduces MSE and MAE by $1.85\%/2.49\%$, $5.56\%/4.12\%$, and $13.66\%/7.58\%$, respectively.
@@
TimeMixer PatchTST Crossformer Leddam ModernTCN MICN TimesNet DLinear
@@
ETTm1 0.344 0.379 0.362 0.391 0.348 0.383 0.355 0.380 0.353 0.382 0.420 0.435 0.354 0.381 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 0.246 0.310 0.269 0.329 0.257 0.319 0.257 0.318 0.256 0.317 0.518 0.501 0.265 0.320 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 0.397 0.424 0.439 0.448 0.404 0.423 0.427 0.441 0.413 0.434 0.440 0.463 0.415 0.430 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 0.341 0.382 0.374 0.406 0.328 0.377 0.349 0.397 0.324 0.381 0.809 0.658 0.345 0.391 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 0.219 0.249 0.233 0.271 0.222 0.262 0.226 0.264 0.226 0.264 0.228 0.287 0.226 0.264 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 0.149 0.245 0.164 0.261 0.161 0.261 0.185 0.284 0.159 0.253 0.181 0.279 0.162 0.256 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 0.360 0.255 0.397 0.282 0.391 0.278 0.409 0.279 0.391 0.264 0.523 0.284 0.452 0.283 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 0.181 0.239 0.200 0.260 0.185 0.254 0.193 0.252 0.194 0.245 0.191 0.242 0.223 0.264 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
@@ 4.2 Short-term Forecasting
Setups. For short-term forecasting, we conduct experiments on the PeMS datasets (Wang et al., 2024b), which capture complex spatiotemporal correlations among multiple variates across city-wide traffic networks. We use mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE) as evaluation metrics. The input length $I$ is set to 96 and the output length $O$ to 12 for all baselines. Details of datasets and metrics are in Appendix D and Sec. C.2.
@@ 4.3 Financial Forecasting
Setups. We conduct experiments on both the U.S. and Chinese stock markets, including the S&P 500 and CSI 500 indices. Stock price movements are influenced by various factors such as economic indicators, market sentiment, geopolitical events, and company-specific news, leading to high non-stationarity. We predict next-day returns using historical data and generate investment portfolios with a buy-hold-sell strategy (Sanderson & Lumpkin-Sowers, 2018). At day $t+1$ open, traders sell day $t$ stocks and buy top-ranked ones based on predicted returns. Following previous work (Lin et al., 2021), we evaluate performance using Annual Return Ratio (ARR), Annual Volatility (AVol), Maximum Drawdown (MDD), Annual Sharpe Ratio (ASR), Calmar Ratio (CR), and Information Ratio (IR). Details of datasets and metrics are in Appendix D and Sec. C.3.
@@ 5 Ablation Studies
Ablation on Integrated and Cointegrated Attention impact and order. We conduct the following experiments: ① Integrated Attention only, ② Cointegrated Attention only, ③ Integrated Attention followed by Cointegrated Attention, and ④ Cointegrated Attention followed by Integrated Attention, with patch downsampling replaced by upsampling in this case. The results in Tab. 5 show that both ① and ② underperform compared to ③, indicating that both components are beneficial. Additionally, ④ shows the weakest performance, possibly because modeling long-term cointegrated relationships first leads to a loss of important short-term temporal features (Wang et al., 2022; Han et al., 2024).
@@ 7 Conclusion
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China, under Grant (62302309, 62171248), Shenzhen Science and Technology Program (JCYJ20220818101014030, JCYJ20220818101012025).
@@ A.2 Cointegration and EG Test
Appendix B Theoretical Analysis of Integrated and Cointegrated Attention
In this section, we provide a more detailed theoretical analysis to justify the design choices behind Integrated Attention and Cointegrated Attention in non-stationary time series modeling. Our approach is grounded in classical stochastic processes, specifically Brownian motion, to explain the rationale behind these mechanisms.
Proposition 1: Spurious Attention from Non-Stationary Inputs
Consider a standard Brownian motion $X_{t}\sim I(1)$, which is a commonly used model for non-stationary processes. We define the process as follows:
$X_{t}=X_{t-1}+u_{t},\quad u_{t}\sim\mathcal{N}(0,\sigma^{2})$
From this, we know that:
$\text{Mean}(X_{t})=0,\quad\text{Var}(X_{t})=t\sigma^{2},\quad\text{Cov}(X_{t_{1}},X_{t_{2}})=\min(t_{1},t_{2})\sigma^{2}$
Let two input patches of length $S$ be:
$p_{i}=[X_{t+i+1},\dots,X_{t+i+S}],\quad p_{j}=[X_{t+j+1},\dots,X_{t+j+S}]$
The attention score between these patches $p_{i}$ and $p_{j}$, based on the dot product between the two vectors, can be approximated as:
$\text{score}(p_{i},p_{j})\propto p_{i}p_{j}^{T}\propto\sum_{s=1}^{S}X_{t+i+s}X_{t+j+s}\propto\sum_{s=1}^{S}(X_{t+i+s}-0)(X_{t+j+s}-0)\propto\sum_{s=1}^{S}\text{Cov}(X_{t+i+s},X_{t+j+s})$
Substituting the covariance expression, we get:
$\text{score}(p_{i},p_{j})\propto\sum_{s=1}^{S}\text{Cov}(X_{t+i+s},X_{t+j+s})\propto\sum_{s=1}^{S}\text{min}(t+i+s,t+j+s)\sigma^{2}\propto\sigma^{2}\left(S\min(i,j)+\frac{S^{2}+2St+S}{2}\right)$
This score grows quadratically with both the time index $t$ and the patch length $S$. As $t$ increases, the score is dominated by long-term trends, leading to spurious attention due to the influence of global trends, rather than capturing genuine short-term dependencies between patches. Figure 8 demonstrates this phenomenon, where many patches show high attention scores due to the accumulation of long-term variance, even though they do not represent meaningful short-term dependencies.
To mitigate this effect, we propose a patch-wise detrending strategy:
$p_{i}^{\prime}=\text{Detrend}(p_{i})=[\Delta X_{t+i},\dots,\Delta X_{t+i+S}]\sim I(0),\quad\Delta X_{t}=X_{t}-X_{t-1}$
This operation removes the non-stationary trend from the input series, leading to a more stable variance:
$\text{Var}(\Delta X_{t})=\sigma^{2}$
Now, the attention score between the detrended patches $p_{i}^{\prime}$ and $p_{j}^{\prime}$ becomes:
$\text{score}(p_{i}^{\prime},p_{j}^{\prime})\propto S\sigma^{2}$
This ensures that the attention mechanism focuses on the genuine short-term dependencies between the patches, unaffected by long-term drift.
Proposition 2: Importance of Non-Stationarity in Capturing Cointegration
Cointegration is a statistical property of a collection of time series variables, where a linear combination of non-stationary series can result in a stationary process. Let us consider two non-stationary time series $X_{t}$ and $Y_{t}$, both following a unit root process $I(1)$:
$X_{t}\sim I(1),\quad Y_{t}\sim I(1)$
If these series are cointegrated, there exists a linear combination of $X_{t}$ and $Y_{t}$ that is stationary:
$Z_{t}=X_{t}-\beta Y_{t}\sim I(0)$
where $\beta$ is a constant coefficient that defines the relationship between the two series.
However, if we remove the non-stationarity by detrending the series, we get:
$\Delta X_{t}=\text{Detrend}(X_{t}),\quad\Delta Y_{t}=\text{Detrend}(Y_{t})$
This leads to the following change in $Z_{t}$:
$Z_{t}=\Delta X_{t}-\beta\Delta Y_{t}=\epsilon_{t}$
where $\epsilon_{t}$ represents a random noise sequence, which destroys the cointegration signal. Removing non-stationary components eliminates the vast majority of cointegration information, as shown in Figure 9.
Thus, to capture long-term dependencies accurately, we must preserve the non-stationarity between the variables. In our framework, Cointegrated Attention (CD) is designed to maintain these long-term dependencies, ensuring that the model can recognize and capture the cointegration between variables across time.
Summary of Propositions
These two propositions explain the necessity of Integrated Attention and Cointegrated Attention:
  • For short-term modeling, we eliminate non-stationarity to avoid spurious regressions and focus on stable, local dependencies within the data.
  • For long-term modeling, we preserve non-stationarity to capture meaningful cointegration relationships between variables, which are crucial for modeling long-term equilibrium dynamics.
Appendix C Metrics
C.1 Long-term Forecasting
@@
C.2 Short-term Forecasting
@@ Appendix D Datasets
Appendix E Implementation Details
All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014). All models are trained for 100 epochs. Following the protocol outlined in the comprehensive benchmark TFB (Qiu et al., 2024), the drop-last trick is disabled during the test phase. Tab. 8 provides detailed hyperparameter settings for each dataset. For the four ETT datasets, the relatively small number of channels results in less pronounced long-term cointegration relationships, as evidenced by the low EG test results in Tab. 7. Therefore, we focus on modeling short-term intra-variate variations only. To improve training stability, we adopt a hybrid MAE loss that operates in both the time and frequency domains. Compared to MSE, which amplifies differences, MAE provides a more stable optimization process. Additionally, the frequency-domain loss mitigates label autocorrelation, making training more effective. The loss function is defined as:
@@ Appendix E Implementation Details
where FFT denotes the Fast Fourier Transform. $\alpha$ is the hyperparameter. We conduct additional experiments where TimeBridge was trained using the MSE loss and compare the results with the best baseline, DeformableTST in Tab. 15. The results show that the hybrid MAE loss improves TimeBridge’s performance on four ETT datasets. The ETT datasets have relatively few channels ($C=7$), which limits TimeBridge ’s ability to utilize the Cointegrated Attention module to capture cointegration information. We only use the Integrated Attention to model ETT datasets (see Tab. 8). Additionally, the ETT data exhibits a high degree of random fluctuation, which is better modeled using a loss function that weighs both time and frequency components. Hence, the hybrid MAE loss strengthens the modeling of short-term dependencies in such datasets. On datasets with more channels (e.g., Electricity, Traffic, Solar), the choice of loss function has minimal impact, as both losses yield similar results.
@@
ETTh1 $3$ 00 $30$ $30$ $2e$-$4$ $128$ $128$ $0.35$
ETTh2 $3$ 00 $15$ $15$ $1e$-$4$ $128$ $128$ $0.35$
ETTm1 $3$ 00 $15$ $15$ $2e$-$4$ $64$ $128$ $0.35$
ETTm2 $3$ 00 $15$ $15$ $2e$-$4$ $64$ $64$ $0.35$
@@
Appendix F Full Results
F.1 Main Experiments